# AutoTAMP: Autoregressive Task and Motion Planning with LLMs as Translators and Checkers

### v1

@@ Abstract
For effective human-robot interaction, robots need to understand, plan, and execute complex, long-horizon tasks described by natural language. The recent and remarkable advances in large language models (LLMs) have shown promise for translating natural language into robot action sequences for complex tasks. However, many existing approaches either translate the natural language directly into robot trajectories, or factor the inference process by decomposing language into task sub-goals, then relying on a motion planner to execute each sub-goal. When complex environmental and temporal constraints are involved, inference over planning tasks must be performed jointly with motion plans using traditional task-and-motion planning (TAMP) algorithms, making such factorization untenable. Rather than using LLMs to directly plan task sub-goals, we instead perform few-shot translation from natural language task descriptions to an intermediary task representation that can then be consumed by a TAMP algorithm to jointly solve the task and motion plan. To improve translation, we automatically detect and correct both syntactic and semantic errors via autoregressive re-prompting, resulting in significant improvements in task completion. We show that our approach outperforms several methods using LLMs as planners in complex task domains. [^1] Datasets and Codes are available at https://github.com/yongchao98/AutoTAMP
Keywords: Large Language Models, Task and Motion Planning, Human-Robot Interaction
1 Introduction
@@
A long-standing goal of robotics has been to provide agents with the ability to find and execute optimal plans for complex tasks. Robots not only need to reason about the task in the environment and find a satisfying sequence of actions but also verify the feasibility of executing those actions according to the robot’s motion capabilities. This problem is referred to as task and motion planning (TAMP), and there has been considerable research on efficient algorithms [1]. Classic solutions rely on specifying tasks in a dedicated planning representation amenable to such algorithms [2, 3, 4, 5].
@@
While this approach to task specification has been quite successful, directly using these representations requires training and experience, making them a poor candidate as an interface for non-expert users. As an alternative, natural language (NL) provides an intuitive and flexible mode of task descriptions. Excitingly, pre-trained large language models (LLMs) have demonstrated surprisingly good performance on many language-related tasks [6], and there has been an associated burst of research investigating their various potential applicability to TAMP [7, 8, 9, 10, 11, 12, 13].
@@
Early efforts showed promising results using LLMs as direct task planners [7], i.e., generating a sequence of sub-tasks based on a set of natural language instructions, but were limited by a lack of feedback and verification of executability of the sub-task sequence. Further research addressed executability by connecting sub-tasks to control policy affordance functions [8], providing environmental feedback of robot actions [10], and interleaving action feasibility checking with LLM action proposals [9]. However, these approaches struggle with various task complexities, such as temporally-dependent multi-step actions, action sequence optimization, task constraints, etc. Furthermore, these frameworks factor the planning problem and use LLMs to infer a task plan separately from the motion plan. In many situations, the task and motion plan must optimized together to fulfill the task. For instance, when the task is ‘reach all locations via the shortest path’, the order of places to be visited (task planning) depends on the geometry of the environment and the related motion optimization. Unfortunately, as we show in this paper, LLMs do not seem capable of directly generating trajectories, possibly related to limitations in complex spatial and numerical reasoning [14, 15].
Notably, the task representations used in classic TAMP, such as PDDL [2] or Temporal Logics [3], are sufficiently expressive to specify task complexities from which existing planning algorithms can find and verify action sequences satisfying those specifications. To benefit from both the user-friendliness of NL and the capabilities of existing TAMP algorithms, we approach the problem by using LLMs to translate from high-level task descriptions to formal task specifications. We are not the first to use LLMs as translators of task specifications [16, 17], but our work contributes several new ideas that improve the performance of such approaches. First, previous work focused on translating natural language into PDDL goals [16] or Linear Temporal Logics (LTL) [18], which only considered the problem of task planning. Here we utilize Signal Temporal Logic (STL) as the intermediary representation, from which the planner directly optimizes the whole trajectory, i.e, task and motion planning together, thus improving planning success rate. Second, we also use re-prompting to automatically correct task specification errors, both via an existing syntactic error correction framework and a novel semantic error checking and correcting loop using LLMs, resulting in significantly improved performance. We also conduct an ablation study integrating a fine-tuned NL-to-STL model [19] with our AutoTAMP approach and show that GPT-4 few-shot learning is competitive with fine-tuning. We conclude that in-context learning with pre-trained LLMs is well suited for language-to-task-specification translation for solving TAMP problems.
In this paper, we contribute the following:
  • Pre-trained LLM (GPT-3 and GPT-4) few-shot translation from NL task descriptions to STLs, and execution of trajectories via an STL planner.
  • A novel autoregressive re-prompting technique for automatic checking and correcting of semantic errors in generated task specifications via LLMs. We show that, when combined with automatic syntactic correction, this technique greatly improves task success rates.
  • Comprehensive experiments in challenging task domains, including several multi-agent tasks. We show that our approach outperforms direct LLM planning methods for tasks with hard geometric and temporal constraints.
  • In addition to our code, we publish a dataset of 1400 test cases consisting of the language instruction, environments, produced STL specifications, and resulting trajectories.
2 Problem Description
Figure 1: GPT-4 failure case for direct end-to-end trajectory planning. The orange line shows the correct path obeying the instruction. The purple and gray dashed lines show the trajectories from GPT-4 after first and second prompts, respectively. GPT-4 generates a list of $(x,y)$ locations with associated timepoints. The initial prompt describes the language modeling task, environment state, and instruction. Each object is a rectangle described by $(x,y)$ boundaries.
@@ 2 Problem Description
3 Methods
Figure 2: Illustration of different approaches applying LLMs for task and motion planning.
@@ 3 Methods
Figure 2 illustrates three of the approaches we compare in our work, each using LLMs in some capacity. Each takes as input (1) a text-based representation of the global environment state, (2) in-context examples for few-shot learning, and (3) a natural language instruction. For methods using feedback or re-prompting, we describe details of the re-prompt context in Appendix A.
3.1 LLM End-to-end Motion Planning
@@
One natural idea is to use an LLM to handle both the task and motion planning by directly generating a trajectory for a given language instruction; we refer to this as LLM End-to-end Motion Planning. Where the generated trajectory violates constraints, we re-prompt the model with the constraint violation to produce another trajectory, allowing up to five such re-prompts. Figure 1 illustrates this pipeline, including a specific failure case with two constraint-violating trajectories.
3.2 LLM Task Planning
@@
A more common approach is to use an LLM to handle the task planning by directly generating a sequence of sub-tasks from a given language instruction; we refer to this as LLM Task Planning. To generate a final trajectory, the sub-tasks are handled by an independent motion planner. In this work, these sub-tasks are limited to navigation actions, and the motion planning is handled by the STL planner used by our proposed approach; this permits fair comparison of results across methods. Each sub-task is converted to STL to be consumed by the planner (more details are presented in Appendix B). We evaluate and compare against three methods that each use LLMs for task planning: (1) Naive Task Planning, (2) SayCan, and (3) LLM Task Planning + Feedback.
@@
Naive Task Planning As proposed by Huang et al. [7], we evaluate using LLMs to generate the entire sub-task sequence without checking for executability.
@@
SayCan Alternatively, an LLM can be autoregressively prompted to generate each subsequent sub-task conditioned on the previous sub-tasks in the sequence. The next sub-task can be selected from the top K candidates by combining the language model likelihood with a feasibility likelihood of the candidate action and choosing the most-likely next sub-task. This is the method proposed by Ahn et al. [8]. We set K to 5 in our evaluations.
@@
LLM Task Planning + Feedback A third task planning method combines full sequence generation with feasibility checking to both find sub-task sequences that satisfy the full task and verify their feasibility before execution. For any infeasible sub-tasks, the LLM can be re-prompted with feedback about the infeasible actions to generate a new sub-task sequence. This is similar to the hierarchical method proposed by Lin et al. [9] but with feedback for re-prompting.
Figure 3: An example of full prompt used by our AutoTAMP framework
3.3 Autoregressive LLM Specification Translation&Checking + Formal Planner
@@
Unlike LLM Task Planning, our approach translates NL to STL with an LLM and then plans the trajectory with an STL planner, as shown in Figure 2. We include two re-prompting techniques to improve translation performance: one for syntactic errors and another for semantic errors. We show one example of the prompt in Figure 3.
@@
where $\phi$ and $\varphi$ are STL formulas, and $\pi^{\mu}$ is an atomic predicate. $\neg$ (negation), $\land$ (and), $\lor$ (or), $\Rightarrow$ (imply), and $\Leftrightarrow$ (equal)) are logical operators. $\textbf{F}_{[a,b]}$ (eventually/finally), $\textbf{G}_{[a,b]}$ (always/globally), and $\textbf{U}_{[a,b]}$ (until) are temporal operators with real-time constraints $t\in[a,b]$. Temporal operators with time constraints are illustrated by Appendix C, and other operators can be presented using the basic syntax. The action primitives in this work are ’enter(room_name)’ and ’not_enter(room_name)’. To avoid issues of parentheses matching, the generated STLs adhere to pre-order form, as discussed in Appendix D.
STL Trajectory Planner We use a state-of-the-art STL planner [21] capable of multi-agent planning. It defines the validity of an STL formula with respect to a trajectory and then optimizes the trajectory to maximize the validity. The planner only considers the state within or out of a region, and outputs a sequence of location & time pairs.
@@
Syntactic Checking & Semantic Checking Open-loop translation can suffer from syntactic and semantic errors. We use two re-prompting techniques to automatically correct such errors. Like Skreta et al. [22] we use a verifier to check for syntax errors; any found are provided as feedback when re-prompting the LLM to generate corrected STL. We repeat until no errors are found (up to five iterations). For semantic errors, we propose a novel autoregressive re-prompting technique that feeds the STL planner’s generated state sequence (i.e., [[in(road), 0], [in(red kitchen), 0.5], [in(blue restroom2), 1.2]…]) back into the LLM to check whether it satisfies the original instruction. If it does not, the LLM is prompted to modify the STL, which repeats the syntactic and semantic re-prompting. This process terminates in the case of no detected errors or no change in STL (up to three iterations).
4 Experimental Design
@@
Each of our task scenarios is set in a 2D environment and entails navigation of one or more robots; the robots have extent in the environment and are initialized with varying start positions. Each environment is comprised of regions with a shape, location, and characteristics (e.g., color, name, function). For each method we evaluate, the LLM is initially prompted with a description of the language task (e.g. task planning) and five in-context examples (these depend on the language task). To mitigate variance across prompts, we initially tested six different sets of examples for each method and chose the one that performed best. Through this testing, we found that the variance over prompts was insignificant relative to overall performance.
@@
We evaluated the different methods described in Section 3 across six different task scenarios (three single-agent and three multi-agent) with different combinations of geometric and temporal constraints. For each scenario description below, we indicate the presence of these constraints below with G and T, respectively. For each method, we evaluate performance with both GPT-3 and GPT-4 as the LLM. Note that in multi-agent scenarios, we do not test SayCan or LLM Task Planning + Feedback because these methods are not straight-forwardly adaptable for multiple agents. We also terminate and report failure for test cases that take more than 90 minutes. We automatically check resulting trajectories via hard-coded checkers. The full set of experiments took two weeks using four 16-core CPUs; the cost of LLM API calls was ~1500USD.
@@
HouseWorld1 (single-agent) As shown in Figure 4(a), this is a house environment from Finucane et al. [23]. We first manually constructed 10 different instructions of varying complexity (see Appendix E) before prompting GPT-4 to paraphrase each into 9 differently worded instructions with the same meaning, resulting in 100 total instructions for this environment. For each instruction, we randomly initialize between two start-end position pairs for 200 total test cases. For this scenario, we do not impose a hard time constraint for the planned trajectory.
Example instruction: ’Visit two rooms with color closest to red, but not the pure red color.’
Corresponding STL: (finally room_purple and finally room_pink).
@@
Example instruction: ’Visit two rooms with color closest to red, but not the pure red color. The task should be completed within 10 seconds.’
Corresponding STL: (finally[0, 10] room_purple and finally[0, 10] room_pink).
@@
The remaining task scenarios were designed with specific rules and goals for the agent(s) to follow. For each scenario, GPT-4 was used to paraphrase the original description into 20 uniquely worded variants with the same meaning. We instantiate three different instances of the environment for each scenario and randomize five different start/end location pairs for a total of 300 test cases. Note that the avoidance of collision among multiple agents has been inherently encoded in the STL planner.
@@
Chip’s Challenge (G, single-agent) Figure 4(b) shows a scenario inspired by a level from Chip’s Challenge, a classic puzzle solving game with strict geometric and logical constraints. The robot must reach all goal regions (blue) but must acquire a unique key to pass through the corresponding door. The task planning must obey both the spatial and logical constraints.
Example instruction: ’Try to reach all the goals but you have to reach the corresponding key first to open the specific door. For example, you have to reach key1 ahead to open door1. Also remember always do not touch the walls.’
[…]
@@
Overcooked (G & T, multi-agent) Figure 4(c) shows a scenario inspired by Overcooked, a popular cooking simulation game with strict time constraints. The agents must cooperatively gather ingredients and return to CookingRoom in a limited time. In this scenario, the multi-agent motion planning is challenged by limited space for agents to maneuver.
Example instruction: ’Enter all the ingredient room to pick up food. Once entered ingredient rooms, go to cooking room within 3 seconds. And all the agents should not collide each other and black obstacles.’
[…]
@@
Rover (G & T, multi-agent) Figure 4(d) is a scenario used by Sun et al. [21]. Multiple agents must reach each observation region (blue goals) before transmitting the observations by entering a red region, all while subjected to time and energy constraints.
Example instruction: ’All rovers must reach the blue charging station within 5 units of time each time they exit it. Once they reach their destination, they need to get to a yellow transmitter within 2 time units to send the collected information to the remote control. Rovers must keep clear of black walls and other rovers. All target areas need to be visited.’
[…]
@@
Wall (G & T, multi-agent) Figure 4(e) is also a scenario from Sun et al. [21]. Multiple agents must occupy each goal region (blue) while subject to a time constraint and a maneuver bottleneck.
Example instruction: ’Reach all the goals and stay there. Take care of the collision among each agent and to the walls.’
[…]
Appendix F shows key frame sequences of correctly planned trajectories for the three multi-agent scenarios as produced via our AutoTAMP method.
5 Results
5.1 Task Success Rates
Table 1: Task success rates in single-agent scenarios. The scenarios have hard temporal or geometric constraints.
HouseWorld1 HouseWorld2 Chip Challenge
(soft time cst.) (hard time cst.) (hard geo. cst.)
GPT-3 end2end $0\%$ $0\%$ $0\%$
GPT-3 naive task planning $74\%$ $36\%$ $0\%$
SayCan $75.5\%$ $36\%$ $0\%$
GPT-3 task planning/feed. 79% $40\%$ $0\%$
GPT-3/STL $28\%$ $27\%$ $29\%$
GPT-3/STL/Syn. $49\%$ $47\%$ $66\%$
GPT-3/STL/Syn./Sem. (AutoTAMP) $62\%$ 62% 74.3%
GPT-4 end2end $9.5\%$ $9.5\%$ $0\%$
GPT-4 naive task planning $90\%$ $45\%$ $0\%$
SayCan $90\%$ $47.5\%$ $0\%$
GPT-4 task planning/feed. 92% $49\%$ $0\%$
GPT-4/STL $43.5\%$ $42\%$ $42.7\%$
GPT-4/STL/Syn. $59.5\%$ $59\%$ $70\%$
GPT-4/STL/Syn./Sem. (AutoTAMP) $82.5\%$ 82% 87.7%

### v3

@@ Abstract
For effective human-robot interaction, robots need to understand, plan, and execute complex, long-horizon tasks described by natural language. Recent advances in large language models (LLMs) have shown promise for translating natural language into robot action sequences for complex tasks. However, existing approaches either translate the natural language directly into robot trajectories or factor the inference process by decomposing language into task sub-goals and relying on a motion planner to execute each sub-goal. When complex environmental and temporal constraints are involved, inference over planning tasks must be performed jointly with motion plans using traditional task-and-motion planning (TAMP) algorithms, making factorization into subgoals untenable. Rather than using LLMs to directly plan task sub-goals, we instead perform few-shot translation from natural language task descriptions to an intermediate task representation that can then be consumed by a TAMP algorithm to jointly solve the task and motion plan. To improve translation, we automatically detect and correct both syntactic and semantic errors via autoregressive re-prompting, resulting in significant improvements in task completion. We show that our approach outperforms several methods using LLMs as planners in complex task domains. See our project website [^4] https://yongchao98.github.io/MIT-REALM-AutoTAMP/ for prompts, videos, and code.
I INTRODUCTION
@@
Providing agents with the ability to find and execute optimal plans for complex tasks is a long-standing goal in robotics. Robots need to not only reason about the task in the environment and find a satisfying sequence of actions but also verify the feasibility of executing those actions given the robot’s motion capabilities. This problem is referred to as task and motion planning (TAMP), and there has been considerable research on efficient algorithms [1]. Classic solutions rely on specifying tasks in a dedicated planning representation, such as PDDL [2] or Temporal logics [3], that is both sufficiently expressive to specify task complexities (e.g. constraints on task execution) and amenable to such algorithms [2, 3, 4, 5].
@@
While this approach to task specification has been quite successful, directly using these representations requires training and experience, making them poor interfaces for non-expert users. As an alternative, natural language (NL) provides an intuitive and flexible way to describe tasks. Pre-trained large language models (LLMs) have demonstrated surprisingly good performance on many language-related tasks [6], and there has been an associated burst of research applying them to task execution [7], task planning [8, 9, 10, 11] and TAMP [12, 13].
@@
Promising early efforts used LLMs as direct task planners [8] generating a sequence of sub-tasks based on a set of natural language instructions, but these approaches were limited by a lack of feedback and inability to verify whether sub-task sequences are executable. Further research addressed executability by connecting sub-tasks to control policy affordance functions [9], providing environmental feedback of robot actions [11], and interleaving action feasibility checking with LLM action proposals [12]; this last work also addressed long-horizon action dependencies. However, these approaches struggle with complex tasks involving temporally-dependent multi-step actions, action sequence optimization [9, 11], and constraints on task execution [12]. Furthermore, these frameworks factor the planning problem and use LLMs to infer a task plan separately from the motion plan. In many situations, the task and motion plan must be optimized together to fulfill the task. For instance, when the task is ‘reach all locations via the shortest path’, the order of places to be visited (task planning) depends on the geometry of the environment and the related motion optimization. Unfortunately, we find that LLMs do not seem capable of directly generating trajectories, possibly due to limitations in complex spatial and numerical reasoning [14, 15].
To benefit from both the user-friendliness of NL and the capabilities of existing TAMP algorithms, we approach the problem by using LLMs to translate from high-level task descriptions to formal task specifications. We are not the first to use LLMs in this way [16, 17], but our work addresses some limitations of prior approaches. Previous work translated natural language to Linear Temporal Logics (LTL) [18], which only considered the problem of task planning, and PDDL problem descriptions [16] or PDDL goals [17]. Here we utilize Signal Temporal Logic (STL) as the intermediary representation, allowing for more expressive constraints than LTL and facilitating integrated task and motion planning as with PDDL [19].
The LLM translation process can produce malformed (syntax errors) and semantically misaligned (semantic errors) formal task specifications. To address syntax errors, we adopt an existing iterative re-prompting technique that relies on an external syntax verifier to prompt the LLM with the specific syntactic error for correction [20]. Unfortunately, the lack of an external verifier makes this technique inapplicable for a semantic misalignment between the original natural language instruction and the translated specification. To address this problem, we contribute a novel autoregressive re-prompting technique that uses an LLM to evaluate whether the generated plan is semantically consistent with the original instruction. We re-prompt the model to check the alignment between the original instruction and the generated plan by providing the context of the instruction, the generated STL, and the output of the planner. We conduct comprehensive experiments in challenging 2D task domains, including several multi-agent tasks, and find that our approach outperforms direct LLM planning for tasks with hard geometric and temporal constraints. We show that, when combined with automatic syntactic correction, our technique significantly improves task success rates. We conduct an ablation study over the translation step by integrating a fine-tuned NL-to-STL model [21] with the AutoTAMP framework and show that GPT-4 few-shot learning is competitive with fine-tuning. In addition to our code, we publish a dataset of 1400 test cases consisting of the language instructions, environments, generated STL, and planner trajectory outputs. We conclude that in-context learning with pre-trained LLMs is well suited for language-to-task-specification translation for solving TAMP problems.
II PROBLEM DESCRIPTION
@@
III METHODS
Figure 1: Illustration of different approaches applying LLMs for task and motion planning; our work contributes the LLM-As-Translator & Checker approach. Each approach accepts a natural language instruction and environment state as input and outputs a robot trajectory.
Figure 2: GPT-4 failure case for direct end-to-end trajectory planning. The orange line shows the correct path obeying the instruction. The purple and gray dashed lines show the trajectories from GPT-4 after first and second prompts, respectively. GPT-4 generates a list of $(x,y)$ locations with associated timestamps. The initial prompt describes the language modeling task, environment state, and instruction. Each object is a rectangle described by $(x,y)$ boundaries.
@@ III METHODS
Figure 1 illustrates three of the approaches we compare in our work, each using LLMs in some capacity. Each takes as input (1) a text-based representation of the global environment state, (2) in-context examples for few-shot learning, and (3) a natural language instruction. The LLM-As-Translator & Checker approach is the contribution of this paper. Details and examples of context for prompting and re-prompting can be found in our code repository4.
III-A LLM End-to-end Motion Planning
@@
One natural idea is to use an LLM for both task and motion planning by directly generating a trajectory for a given language instruction; we refer to this as LLM End-to-end Motion Planning. In cases where the generated trajectory violates constraints, we re-prompt the model with the constraint violation to produce another trajectory, allowing up to five such re-prompts. Figure 2 shows this pipeline, including a specific failure case with two constraint-violating trajectories. The LLM End-to-end Motion Planning violates the constraints multiple times even with direct correction re-prompts. This results from the poor spatial and numerical reasoning abilities of LLMs.
III-B LLM Task Planning
@@
A more common approach is to use an LLM to handle the task planning by directly generating a sequence of sub-tasks from a given language instruction; we refer to this as LLM Task Planning. To generate a final trajectory, the sub-tasks are handled by an independent motion planner. In this work, these sub-tasks are limited to navigation actions, and the motion planning is handled by the STL planner used by our proposed approach; this permits fair comparison of results across methods. Each sub-task is converted to STL to be consumed by the planner. We evaluate and compare against three methods that each use LLMs for task planning: (1) Naive Task Planning, (2) SayCan, and (3) LLM Task Planning + Feedback.
@@
Naive Task Planning As proposed by [8], we evaluate using LLMs to generate the entire sub-task sequence without checking for executability.
@@
SayCan Alternatively, an LLM can be iteratively prompted to generate each subsequent sub-task conditioned on the previous sub-tasks in the sequence. The next sub-task can be selected from the top K candidates by combining the language model likelihood with a feasibility likelihood of the candidate action and choosing the most-likely next sub-task, as proposed by [9]. We set K to 5 in our evaluations.
@@
LLM Task Planning + Feedback A third task planning method combines full sequence generation with feasibility checking to both find sub-task sequences that satisfy the full task and verify their feasibility before execution. For any infeasible sub-tasks, the LLM can be re-prompted with feedback about the infeasible actions to generate a new sub-task sequence. This is similar to the hierarchical method proposed by [12] but with feedback for re-prompting.
III-C Autoregressive LLM Specification Translation&Checking + Formal Planner
@@
Unlike LLM Task Planning, our approach translates NL to STL with an LLM and then plans the trajectory with an STL planner, as shown in Figure 1. We include two re-prompting techniques to improve translation performance: one for syntactic errors and another for semantic errors. By “semantic error”, we mean a misalignment between the intended task described in natural language and the STL expression to which it is translated. Figure 3 shows the structure of the context for re-prompting the model for semantic error correction; we include a full prompt example in our code repository4.
@@
where $\phi$ and $\varphi$ are STL formulas, and $\pi^{\mu}$ is an atomic predicate. $\neg$ (negation), $\land$ (and), $\lor$ (or), $\Rightarrow$ (imply), and $\Leftrightarrow$ (equal)) are logical operators. $\textbf{F}_{[a,b]}$ (eventually/finally), $\textbf{G}_{[a,b]}$ (always/globally), and $\textbf{U}_{[a,b]}$ (until) are temporal operators with real-time constraints $t\in[a,b]$. The action primitives in this work are ’enter(room_name)’ and ’not_enter(room_name)’.
STL Trajectory Planner We use a state-of-the-art multi-agent STL planner [23] that uses piece-wise linear reference paths defined by timed waypoints to recursively encode the constraints expressed in the provided STL expression. It defines the validity of an STL formula with respect to a trajectory and then optimizes the trajectory to maximize the validity. The planner not only searches for a sub-task sequence but also optimizes the time efficiency under dynamical constraints of robot maximum velocity. Here we assume that the locations and shapes of all the objects/rooms in the whole environment are known, which serves as the environment information to the STL planner.
@@
Syntactic Checking & Semantic Checking Open-loop translation can suffer from syntactic and semantic errors. We use two re-prompting techniques to automatically correct such errors. Like [20], we use a verifier to check for syntax errors (we use a simple rules-based STL syntax checker); any errors are provided as feedback when re-prompting the LLM to generate corrected STL. We repeat until no errors are found (up to five iterations). For semantic errors, we propose a novel autoregressive re-prompting technique; we provide the STL planner’s generated state sequence (i.e., [[in(road), 0], [in(red kitchen), 0.5], [in(blue restroom2), 1.2],…]) as context alongside the original instruction and ask the LLM to check whether the plan aligns with the instruction’s semantics. If it does not, the LLM is prompted to modify the STL, which repeats the syntactic and semantic re-prompting. This process terminates in the case of no detected errors or no change in STL (up to three iterations). The structure of the semantic error prompt is shown in Figure 3; full example prompts can be found in our code repository4.
Figure 3: High-level structure of the prompt used for AutoTAMP. The arrow on the right indicates re-prompting for syntax error correction. The arrow on the left indicates re-prompting in cases of semantic errors.
IV EXPERIMENTAL DESIGN
@@
Each task scenario is set in a 2D environment and entails navigation of one or more robots; the robots have extent in the environment and are initialized with varying start positions. Each environment consists of regions with shapes, locations, and properties (e.g., color, name, function). For each method, the LLM is initially prompted with a description of the language task (e.g. task planning or translation) and five in-context examples for that task. To mitigate variance across prompts, we initially tested six different sets of examples for each method and chose the one that performed best. Through this testing, we found that the variance over prompts was insignificant relative to overall performance.
@@
We evaluated the different methods described in Section III across six different task scenarios (three single-agent and three multi-agent) with different combinations of geometric and temporal constraints. For each scenario description below, we indicate the presence of these constraints below with G and T, respectively. For each method, we evaluate performance with both GPT-3 and GPT-4 as the LLM. Note that in multi-agent scenarios, we do not test SayCan or LLM Task Planning + Feedback because these methods are not straight-forwardly adaptable for multiple agents. For multi-agent tasks, the agents are assigned a subtask and a time for completion at each time step; since the time for completion is often different, it is not obvious how/when to check and provide feedback. We also terminate and report failure for test cases that take more than 90 minutes. We automatically check resulting trajectories via hard-coded checkers. The full set of experiments took two weeks using four 16-core CPUs; the cost of LLM API calls for evaluating all of the approaches was $\sim$1500 USD.
@@
HouseWorld1 (single-agent) As shown in Figure 4(a), this is a house environment from [24]. We first manually constructed 10 different instructions of varying complexity before prompting GPT-4 to paraphrase each into 9 differently worded instructions with the same meaning, resulting in 100 total instructions for this environment. For each instruction, we randomly initialize between two start-end position pairs for 200 total test cases. For this scenario, we do not impose a hard time constraint for the planned trajectory.
@@
The remaining task scenarios were designed with specific rules and goals for the agent(s) to follow. For each scenario, GPT-4 was used to paraphrase the original description into 20 unique variants with the same meaning, which are further checked by humans. We instantiate three different instances of the environment for each scenario and randomize five different start/end location pairs for a total of 300 test cases.
@@
Chip’s Challenge (G, single-agent) Figure 4(b) shows a scenario inspired by Chip’s Challenge, a classic puzzle solving game with strict geometric and logical constraints. The robot must reach all goal regions (blue) but must acquire a unique key to pass through the corresponding door.
@@
Overcooked (G & T, multi-agent) Figure 4(c) shows a scenario inspired by Overcooked, a popular cooking simulation game with strict time constraints. The agents must cooperatively gather ingredients and return to CookingRoom in a limited time. The multi-agent motion planning is challenged by limited space for agents to maneuver.
@@
Rover (G & T, multi-agent) Figure 4(d) is a scenario used by [23]. Multiple agents must reach each observation region (blue) before transmitting their observations from a red region, all while subjected to time and energy constraints.
@@
Wall (G & T, multi-agent) Figure 4(e) is also from [23]. Multiple agents must occupy each goal region (blue) while subject to a time constraint and a maneuver bottleneck.
V RESULTS